OpenStudy (anonymous):

Compute the flux of the vector field, vector F= 4x^3i + 9xyj + 9xzk, through the surface shown below. The surface is a cylinder with radius 1 and length 2, oriented away from the x-axis. 19

OpenStudy (anonymous):

help me @sidsiddhartha

OpenStudy (anonymous):

please :)

OpenStudy (anonymous):

help me please @perl @ganeshie8

OpenStudy (anonymous):

@eliassaad can you also help me with this one please

OpenStudy (anonymous):

@eliassaab

OpenStudy (anonymous):

Does the surface includes top and bottom?

OpenStudy (anonymous):

Use Firefox on my site http://moltest.missouri.edu/mucgi-bin/calculus.cgi and choose Calculus III (Vector Calculus)

OpenStudy (anonymous):

no the surface does not includes top or bottom

OpenStudy (anonymous):

Is the answer 0?

OpenStudy (anonymous):

Let us divide the surface into two \(S_1 \) above the x y plane and \(S_2 \) below the xy plane

OpenStudy (anonymous):

I got 36pi

OpenStudy (anonymous):

\[\int\limits_{0}^{2\pi}\int\limits_{0}^{2}(4x^3i+9xcos \theta j+9x \sin \theta k) * (\cos \theta j + \sin \theta k) dxd \theta \]

OpenStudy (anonymous):

and that gives me 36pi

OpenStudy (anonymous):

@eliassaab

OpenStudy (anonymous):

36pi is right :)